OpenStudy (anonymous):

Oscar bought a bag of almonds. He ate 3/8 of the bag on Sunday. On Monday he ate 2/3 of the almonds left. What fraction of the entire bag did he eat on Monday?

OpenStudy (anonymous):

let x be the number of almonds he ate 3/8 so x-(3/8)x=x(1-3/8) left monday x(5/8)-2/3x

OpenStudy (texaschic101):

He ate 3/8 of the bag leaving 5/8. (Sunday) 2/3 of 5/8 (Monday) 2/3 * 5/8 = (just multiply and you will have your answer) You might need a second opinion on this because I am not sure

OpenStudy (anonymous):

@texaschic101 is correct the is differnece between eating a fraction of the bag and a fraction of the almonds

OpenStudy (texaschic101):

so would you multiply 2/3 by 5/8 or would you subtract 2/3 from 5/8 ???

OpenStudy (texaschic101):

@Jonask ....???

OpenStudy (anonymous):

multiply

OpenStudy (texaschic101):

thanks...I got confused for a second.

OpenStudy (anonymous):

is you try subtracting you actually get a negetive answer????!!!!